I need help on a problem. It is to solve a system of equations by the substitution method. The system is:
x+3y-1
x+2y=9

Answer by elima(1433)

x+3y-1
x+2y=9
First solve one of the equations for x;
x+3y=1
x=1-3y
Now substitute that into the other equation for x;
x+2y=9
(1-3y)+2y=9
1-y=9
-y=9-1
-y=8
y=-8
Now plug the answer for y into the equation to get x;
x+2(-8)=9
x-16=9
x=9+16
x=25
Hope you understand
=)